OpenStudy (lukecrayonz):

For f(x)=-4x^2-x-1, find f(x+h)-f(x)/h I solved it but I don't know exactly how to fill it out? I'm using MyMathLab

OpenStudy (psymon):

MyMathLab is annoying -_- Maybe it's just being REALLY picky about how you input it?

OpenStudy (lukecrayonz):

I agree its annoying haha, but yeah, I don't know how to put it in as they want it.

OpenStudy (psymon):

You got the answer on ya real quick?

OpenStudy (lukecrayonz):

i THINK its (-8hx-4h^2-2x-h-2)/h

OpenStudy (psymon):

Not quite. This is what you would have if done properly: \[\frac{ -8xh-4h ^{2}-h }{ h }\]But even if you have this part correct, you have more workto do. See if you can figure out how I got that. If not then I can work ya through it.

OpenStudy (lukecrayonz):

I understand :-). It's really late here ( 6 a.m) So then for dividing h, my answer is -8x-4h?

OpenStudy (psymon):

Almost. watch what happens when I factor h out of the top: \[\frac{ -8xh }{ h }-\frac{ 4h ^{2} }{ h }-\frac{ h }{ h }=-8x-4h-1 \] see why there's a 1 there still?

OpenStudy (lukecrayonz):

Ahh right dividing h/h, forgot. One more question!

OpenStudy (lukecrayonz):

Is this correct? http://gyazo.com/8da7fd0a6abfc2450ff3d5f910a46ba7

OpenStudy (psymon):

Sure

OpenStudy (lukecrayonz):

I'll post to another website if you don't like the look of that one :0 I get that a lot.

OpenStudy (psymon):

Nope. domain is how much of the x-values the graph covers. The line doesn't go all the way to 100, so it can't be that.

OpenStudy (lukecrayonz):

Thank you! I got a 100:-). Time for sleep! I will need your help some other time!

OpenStudy (psymon):

Lol, enjoy the sleep then. Good job :3